# Time-domain study of coupled collective excitations in quantum materials

### v1

@@ Acknowledgements.
We thank helpful input from Edoardo Baldini, Henry Bell, Thomas Blank, Yu-Che Chien, Shaofeng Duan, Nuh Gedik, Patrick Liu, Tianchuang Luo, Mariano Trigo, and Xunqing Yin during the preparation of this manuscript.References
@@
- Ilyas et al. [2024] B. Ilyas, T. Luo, A. von Hoegen, E. Viñas Boström, Z. Zhang, J. Park, J. Kim, J.-G. Park, K. A. Nelson, A. Rubio, and N. Gedik, Terahertz field-induced metastable magnetization near criticality in FePS3, Nature 636, 609 (2024).
@@
- Robert et al. [2016a] C. Robert, D. Lagarde, F. Cadiz, G. Wang, B. Lassagne, T. Amand, A. Balocchi, P. Renucci, S. Tongay, B. Urbaszek, and X. Marie, Exciton radiative lifetime in transition metal dichalcogenide monolayers, Phys. Rev. B 93, 205423 (2016a).
@@
- Zong et al. [2021] A. Zong, P. E. Dolgirev, A. Kogar, Y. Su, X. Shen, J. A. W. Straquadine, X. Wang, D. Luo, M. E. Kozina, A. H. Reid, R. Li, J. Yang, S. P. Weathersby, S. Park, E. J. Sie, P. Jarillo-Herrero, I. R. Fisher, X. Wang, E. Demler, and N. Gedik, Role of equilibrium fluctuations in light-induced order, Physical Review Letters 127, 227401 (2021).
@@
- Sheng et al. [2024] S. Sheng, M. Abdo, S. Rolf-Pissarczyk, K. Lichtenberg, S. Baumann, J. A. J. Burgess, L. Malavolti, and S. Loth, Terahertz spectroscopy of collective charge density wave dynamics at the atomic scale, Nat. Phys. 10.1038/s41567-024-02552-7 (2024).
@@
- Robert et al. [2016b] C. Robert, D. Lagarde, F. Cadiz, G. Wang, B. Lassagne, T. Amand, A. Balocchi, P. Renucci, S. Tongay, B. Urbaszek, and X. Marie, Exciton radiative lifetime in transition metal dichalcogenide monolayers, Phys. Rev. B 93, 205423 (2016b).
@@
- Blank et al. [2023] T. G. H. Blank, K. A. Grishunin, B. A. Ivanov, E. A. Mashkovich, D. Afanasiev, and A. V. Kimel, Empowering control of antiferromagnets by THz-induced spin coherence, Physical Review Letters 131, 096701 (2023).

### v2

@@ Acknowledgements.
We thank helpful input from Edoardo Baldini, Henry Bell, Thomas Blank, Yu-Che Chien, Shaofeng Duan, Nuh Gedik, Patrick Liu, Megan Loh, Tianchuang Luo, Enrico Perfetto, Ruixin Qiu, Gianluca Stefanucci, Mariano Trigo, and Xunqing Yin during the preparation of this manuscript.References
@@
- Ilyas et al. [2024] B. Ilyas, T. Luo, A. von Hoegen, E. Viñas Boström, Z. Zhang, J. Park, J. Kim, J.-G. Park, K. A. Nelson, A. Rubio, and N. Gedik, Terahertz field-induced metastable magnetization near criticality in FePS3, Nature 636, 609–614 (2024).
@@
- Robert et al. [2016] C. Robert, D. Lagarde, F. Cadiz, G. Wang, B. Lassagne, T. Amand, A. Balocchi, P. Renucci, S. Tongay, B. Urbaszek, and X. Marie, Exciton radiative lifetime in transition metal dichalcogenide monolayers, Phys. Rev. B 93, 205423 (2016).
@@
- Zong et al. [2021] A. Zong, P. E. Dolgirev, A. Kogar, Y. Su, X. Shen, J. A. W. Straquadine, X. Wang, D. Luo, M. E. Kozina, A. H. Reid, R. Li, J. Yang, S. P. Weathersby, S. Park, E. J. Sie, P. Jarillo-Herrero, I. R. Fisher, X. Wang, E. Demler, and N. Gedik, Role of equilibrium fluctuations in light-induced order, Phys. Rev. Lett. 127, 227401 (2021).
@@
- Perfetto et al. [2016] E. Perfetto, D. Sangalli, A. Marini, and G. Stefanucci, First-principles approach to excitons in time-resolved and angle-resolved photoemission spectra, Phys. Rev. B 94, 245303 (2016).
- Steinhoff et al. [2017] A. Steinhoff, M. Florian, M. Rösner, G. Schönhoff, T. O. Wehling, and F. Jahnke, Exciton fission in monolayer transition metal dichalcogenide semiconductors, Nat. Commun. 8, 1166 (2017).
@@
- Rustagi and Kemper [2019] A. Rustagi and A. F. Kemper, Coherent excitonic quantum beats in time-resolved photoemission measurements, Phys. Rev. B 99, 125303 (2019).
- Christiansen et al. [2019] D. Christiansen, M. Selig, E. Malic, R. Ernstorfer, and A. Knorr, Theory of exciton dynamics in time-resolved ARPES: Intra- and intervalley scattering in two-dimensional semiconductors, Phys. Rev. B 100, 205401 (2019).
@@
- Sheng et al. [2024] S. Sheng, M. Abdo, S. Rolf-Pissarczyk, K. Lichtenberg, S. Baumann, J. A. J. Burgess, L. Malavolti, and S. Loth, Terahertz spectroscopy of collective charge density wave dynamics at the atomic scale, Nat. Phys. 20, 1603–1608 (2024).
@@
- Blank et al. [2023] T. G. H. Blank, K. A. Grishunin, B. A. Ivanov, E. A. Mashkovich, D. Afanasiev, and A. V. Kimel, Empowering control of antiferromagnets by THz-induced spin coherence, Phys. Rev. Lett. 131, 096701 (2023).
@@
- Zhang et al. [2024d] Z. Zhang, Y.-C. Chien, M. Tou Wong, F. Y. Gao, Z.-J. Liu, X. Ma, S. Cao, E. Baldini, and K. A. Nelson, Terahertz stimulated parametric downconversion of a magnon mode in an antiferromagnet (2024d), arXiv:2412.01989 .
@@
- Baum and Ropers [2024] P. Baum and C. Ropers, Comment on ’Attosecond electron microscopy and diffraction’ (2024), arXiv:2411.14518 .